# Fine-Grained Assertion-Based Test Selection

### v1

@@ III-D Test Selection
Test class level selects all test methods in a test class. Test classes, with special features (as described in Section III-C), that are parameterized, use inheritance, or contain test cases calling other tests, are selected at the test class level for the sake of safety and efficiency. Algorithm 2 summarizes the procedure for our test selection. With the test dependencies $\mathcal{D}$ obtained by running the instrumented subject, the assertion slices $\mathcal{A}$ and the changes $\delta$ computed by Algorithm 1, the selected test entities $\Gamma$ for the current code revision can be computed by Algorithm 2. $\Gamma_{A}$, $\Gamma_{M}$, $\Gamma_{C}$ denote selected assertion slices, test methods and test classes, respectively. We categorize the code changes obtained from Algorithm 1 into two distinct sets: test code changes, denoted as $\delta_{T}$, and production code changes, denoted as $\delta_{P}$. For each changed test entity $t\in\delta_{T}$ which could be a modified or newly added test method or test class, its assertion slices need to be updated to align with the current code reversion (Line 2-3). Additionally, $\delta_{T}$ should also be selected for execution (Line 25). For each production code change $p\in\delta_{P}$, Line 6 computes the test entities $\mathcal{T}_{p}$ that depend on the production code change $p$ based on the dependency files in $\mathcal{D}$. Because there are some special cases that test dependencies are not collected at the test statement level (see Section III-C), the returned test entities $\mathcal{T}_{p}$ might contain test methods and test classes. For each affected test entity $t\in\mathcal{T}_{p}$, if $t$ is a test statement (Line 9), we retrieve the test method $m$ that contains $t$ (Line 10) and check if there are assertion slices available for the test method $m$ (Line 11). If assertion slices exist for $m$, all the assertion slices that contain $t$ are selected (Line 11-14). If there are no assertion slices for $m$, our selection is performed at the test method level (Line 16-17) to guarantee a safe selection. If test entity $t$ is not a test statement, $t$ is selected as the test method or test class instead (Line 20). Last but not least, all the selected test entities $\Gamma_{A}$, $\Gamma_{M}$, $\Gamma_{C}$, $\delta_{T}$ are used to rewrite their corresponding test classes to make sure only the selected assertions, methods or classes will be executed (Line 25). After the execution is finished, the test classes are restored to the original status, while the rewritten version is saved as a reference.
@@ IV Evaluation
  • RQ4 (performance): What is the performance of Selertion in terms of analysis overhead and collection time? Is there any performance difference when disabling the assertion-level selection?
@@ IV-A Existing RTS Tools
Our search criteria for existing test selection techniques to compare with included tools that are available, runnable, and target Java/JUnit. We selected two state-of-the-art tools, namely HyRTS [12] and Ekstazi [9]. To the best of our knowledge, these are the exclusive options that currently meet our criteria. Significantly, they outperform older RTS techniques like FaultTracer [11], which exhibit slower performance and offer test method selection, as documented in prior research [9]. HyRTS conducts the analysis of test dependencies at either the method or class level and selects test entities at the test class level. Ekstazi analyzes production code to identify changes at the class level. Although the original paper [9] mentions the provision of both method and class-level selection, the Ekstazi tool [23] is currently available with default settings that primarily select tests at the test class level. Our technique collects test dependencies also at a mix of method and class level similar to HyRTS, however, it has the ability to select test entities at multiple levels from the coarse-grained class to the fine-grained assertion level based on its analysis of the test suite. Therefore, theoretically speaking, Selertion should be able to select tests more precisely than both HyRTS and Ekstazi. We empirically assess this in our evaluation.
@@ IV-B Subject Systems
We include a total of 11 subject systems. These subjects are selected based on their relevance to prior research on test assertion analysis [34, 14] and regression test selection [12] and their popularity on GitHub. To ensure fair comparison among Selertion, HyRTS and Ekstazi, our selection criteria include subjects that use Maven with executable and passing JUnit4 test cases. We focus on subjects with a test runtime of at least 30 seconds. As reported by others [12, 9], for lower test runtimes it can take more time to analyze the code than to execute the entire test suite. The selected subjects and their characteristics are shown in Table I with the ascending order of their testing time. All the subjects are single-module projects, except Tika [30] and Accumulo [31]. Because some modules of these multi-module projects only have a few test cases, for simplicity, we choose the module with most test cases to analyze (i.e., tika-parsers module for Tika, core module for Accumulo). We examine the test suites for each subject and blacklist test classes that do not meet the selection criteria (e.g., test classes that mix JUnit3 and JUnit4 APIs in OpenTripPlanner). As some of the subjects are migrated to JUnit5 [35] recently, for consistency, we start from a revision that uses JUnit4 as the Head to collect 100 revisions before it, and remove revisions that fail to build or have no code changes. For the three subjects commons-exec, commons-net and commons-pool, which lack a sufficient number of revisions with substantial source code changes, we introduce changes artificially. To achieve this, we use a source code mutation generator called Major [36] to create random mutants in various production methods, thus introducing more prevalent source code changes. The selected subjects span different application domains and exhibit diverse code and test suite sizes. Table I presents the short hash under the Head column, and lines of code for both production and test code for the Head revision. We measure lines of code using Cloc [37]. The column Time details the average time required to execute all the tests using the build command (i.e., mvn test) for the revisions we analyzed, along with the average execution time for each test case. This table also presents information pertaining to test cases and test assertions that are further elaborated in the following subsection.
@@ IV-C Procedure and Results
Given that we exclude parameterized tests, inherited tests, and tests dependent on other tests from our assertion-level selection (such tests are selected at method or class level), with consideration for the safety of our analysis, it is important to know to what extent the test suite for each subject can be fine-grained selected at the assertion-level. To measure the prevalence of fine-grained selection, we implemented a tool using JDT [19]. This tool statically counts the total number of test cases that can be fine-grained selected for every revision we executed. We present the average values for each subject in Table I. Additionally, because the smallest unit of interest for our study is the test assertion, we count the total number of assertions and calculate the average number per test. These numbers are included under the Assertions column in the table, both in total and per test. It’s worth noting that we opt to count these numbers statically instead of relying on information obtained from executed tests (e.g., as provided by maven-surefire-plugin [38]) to avoid repetitive counting, as a single test case may be executed multiple times due to inheritance or different parameters. The column Fine-grained in Table I shows the number and percentage of test cases that can be fine-grained selected by Selertion in the subject systems. The results show on average 48% out of the total 8,328 test cases can be selected at the assertion level. The percentages range from 11% (commons-pool) to 75% (asterisk-java). Moreover, if each test case contains less than two assertions, the benefits of fine-grained selection become limited because the assertion-level selection will be equivalent to the method-level selection (i.e., the whole test method is selected). Therefore, we count explicit assertion statements (e.g., assertTrue ) inside each test method body. Column Assertions in Table I shows the number of assertions for each subject, which totals 28,177 for all subjects. The average number of assertions per test case ranges from 2.1 for commons-exec to 4.8 for tika-parsers, and the overall average of all the subjects is 3.3 assertions per test case. This indicates that (1) most test cases contain multiple assertions in practice, (2) tests that are analyzable by our approach at the fine-grain assertion-level are prevalent, and (3) our fine-grained selection approach is feasible to be applied to real-world test suites.
@@
To assess the efficacy of our approach in terms of precision, we use the widely used RTS metric Selected Test Ratio [39, 9, 40, 41, 10, 11, 12] to compare with the other two RTS tools. Selected Test Ratio is defined as the ratio of the number of selected test cases to the number of total test cases. As the fine-grained selection is at the test assertion level in this study, we define a new metric Selected Assertion Ratio as the ratio of the number of selected assertions to the number of total assertions. Lower selected test/assertion ratio indicates better selection precision, as it signifies the selection of fewer tests. Our precision evaluation results are presented in Table II and Figure 4. In Table II, the macro-columns Selected Assertion Ratio and Selected Test Ratio present the mean values of all the revisions we studied for Selertion, HyRTS and Ekstazi, respectively, except for commons-exec, tabula-java and tika-parsers. In the case of these subjects, HyRTS failed to run due to StackOverflow errors. Therefore, we mark the failed subjects and average values for HyRTS with an asterisk to signify that the values are calculated with subjects excluding the ones that failed, leading to average calculation different from those of Selertion and Ekstazi. As our results show, Selertion has lower selected test ratio values, which means it selects fewer test cases and assertions than the other two tools. Regarding selected assertion ratio, on average, Selertion selects 15.10% out of all the test assertions and all subjects, which is 16.42% less than Ekstazi. When compared with HyRTS (excluding the marked subjects), Selertion is 7.34% less than HyRTS’s average of 17.96%. For selected test ratio, overall our tool selects 15.80% of all the test cases (10.85% when excluding the marked subjects). This is 8.31% less than HyRTS (excluding the marked subjects) and 17.78% less than Ekstazi. For each subject, the precision increase in terms of selected assertion ratio is from 0.7% (OpenTripPlanner) to 22.3% (LogicNG) in comparison with HyRTS, and from 4.6% (OpenTripPlanner) to 42.9% (LogicNG) compared with Ekstazi. Meanwhile, our tool decreases the selected test ratio from 3.9% (OpenTripPlanner) to 22.5% (LogicNG) in comparison to HyRTS, and from 7.7% (accumulo-core) to 44.5% (LogicNG) compared with Ekstazi. It is worth mentioning that if the precision improvement is smaller for one subject than the other, it does not mean the subject performs worse or vice versa. Rather, the selected ratios are related to the size of code changes in each sample revision we collect to analyze. Larger code changes always lead to higher selected ratios as well as higher ratio differences between tools (e.g., LogicNG). Moreover, to provide more insights about selected assertion ratios for each subject besides the mean values, Figure 4 includes boxplots depicting the distributions of selected assertion ratios for each revision. The mean values are marked as diamonds, and median values are marked as lines. The results indicate that our approach has the ability to select tests more precisely than other coarser-grained tools.
@@
Because we choose to analyze source code instead of bytecode in our implementation, to assess whether Selertion is efficient to reduce testing time, we measure the offline End-to-End time (AE time) which includes the analysis (A) time to compute code changes, select affected test entities and update assertion slices, as well as the execution (E) time for the selected test entities, then normalize the end-to-end time into a ratio with respect to the overall time of executing the whole test suite (RetestAll). We perform the measurement for each revision and report the average. For fair comparison, we also measure the offline end-to-end time and its ratio for HyRTS and Ekstazi respectively. The results are shown in Table III. Column RetestAll shows the time required to run the whole test suite in average for each subject as a reference. The other three macro-columns present the offline end-to-end time and its ratio to RetestAll for Selertion, HyRTS, and Ekstazi separately. Besides the average values in Table III, to further understand the results, we perform Paired T-Test [42] with the significance level as 5% ($\alpha$ = 0.05) to determine if there is a significant difference between each two groups (i.e., Selertion vs. HyRTS, Selertion vs. Ekstazi and HyRTS vs. Ekstazi) for all the revisions of every subject. The reason to choose Paired T-Test is that as a parametric test, it has greater statistical power than non-parametric tests like Wilcoxon Test [42], and can provide trustworthy results even for non-normal distributions. We highlight the best results (shortest running time) in yellow in each row across the three tools. If the end-to-end time has no statistical difference in comparison with others, it means it has competitive or equivalent capability for reducing test time. After examining the results of offline end-to-end time, in Table III, we mark the subject as light gray if Selertion performs competitively or better than one of the other tools; otherwise, Selertion outperforms the other two tools for that subject. Across all 11 subjects, Selertion consistently reduces testing time effectively, with ratios to RetestAll ranging from 14.9% (commons-net) to 70.3% (OpenTripPlanner). This reduction translates to an overall decrease in testing time of 29.7% to 85.1%. For asterisk-java with 35 seconds runtime, our tool is competitive with HyRTS, but slower than Ekstazi. For subject commons-net, although HyRTS has the shortest running time, Selertion is only 1 second slower than HyRTS and is equivalent to Ekstazi, which has no statistical difference in comparison with both tools. For subject commons-math, Selertion outperforms Ekstazi but is 5.3 seconds slower than HyRTS.
@@
Longer Test Runtime. For the subjects with runtimes longer than 86 seconds, Selertion performs significantly better than Ekstazi in terms of testing time. HyRTS fails to run on the subjects commons-exec, tabula-java and tika-parsers, thus we treat Selertion as superior on these three subjects. As an exception, commons-math has long test time with 160 seconds, the reason that HyRTS works better than Selertion is the execution time per test for commons-math is very fast with only 0.03s (shown in Table I). Therefore, although test assertions are reduced significantly, the reduced execution time is marginal. Although HyRTS has lower average testing time for commons-math (row 8), the result of Paired T-Test indicates that there is no significant difference between Selertion and HyRTS. Selertion performs better than HyRTS for the remaining subjects. Generally speaking, we can conclude that Ekstazi works best for subjects with short test runtime such as asterisk-java with 35s running time, Selertion outperforms or is competitive to both HyRTS and Ekstazi with long-running subjects i.e., test runtime longer than 86 seconds in our experiments. The results show that our fine-grained RTS technique is more effective in handling time-intensive tests. Considering all test suites in our subjects, short and long, the total end-to-end time on average for Selertion is 74 seconds which is 10 second less than HyRTS and 46 seconds less than Ekstazi. The average gain compared with HyRTS could be potentially larger if HyRTS were executable on commons-exec, tabula-java and tika-parsers. To the ratio of RetestAll, our technique reduces 63% of the overall test execution time, which is 5% more than HyRTS (again if we ignore commons-exec, tabula-java and tika-parsers) and significantly more than Ekstazi. Our results show that Selertion is more precise and more efficient, especially for test suites with longer execution times, which benefit most from test selection in practice.
@@
Ave. 200.36 74.33 - 83.51* - 119.74 -
IV-C4 Performance (RQ4).
TABLE IV: Performance results for Selertion and comparison to a method-level variant

### v2

@@ III-D Test Selection
Test class level selects all test methods in a test class. Test classes, with special features (as described in Section III-C), that are parameterized, use inheritance, or contain test cases calling other tests, are selected at the test class level for the sake of safety and efficiency.
Algorithm 2 summarizes the procedure for our test selection. With the test dependencies $\mathcal{D}$ obtained by running the instrumented subject, the assertion slices $\mathcal{A}$ and the changes $\delta$ computed by Algorithm 1, the selected test entities $\Gamma$ for the current code revision can be computed by Algorithm 2. $\Gamma_{A}$, $\Gamma_{M}$, $\Gamma_{C}$ denote selected assertion slices, test methods and test classes, respectively. We categorize the code changes obtained from Algorithm 1 into two distinct sets: test code changes, denoted as $\delta_{T}$, and production code changes, denoted as $\delta_{P}$. For each changed test entity $t\in\delta_{T}$ which could be a modified or newly added test method or test class, its assertion slices need to be updated to align with the current code reversion (Line 2-3). Additionally, $\delta_{T}$ should also be selected for execution (Line 25). For each production code change $p\in\delta_{P}$, Line 6 computes the test entities $\mathcal{T}_{p}$ that depend on the production code change $p$ based on the dependency files in $\mathcal{D}$. Because there are some special cases that test dependencies are not collected at the test statement level (see Section III-C), the returned test entities $\mathcal{T}_{p}$ might contain test methods and test classes. For each affected test entity $t\in\mathcal{T}_{p}$, if $t$ is a test statement (Line 9), we retrieve the test method $m$ that contains $t$ (Line 10) and check if there are assertion slices available for the test method $m$ (Line 11). If assertion slices exist for $m$, all the assertion slices that contain $t$ are selected (Line 11-14). If there are no assertion slices for $m$, our selection is performed at the test method level (Line 16-17) to guarantee a safe selection. If test entity $t$ is not a test statement, $t$ is selected as the test method or test class instead (Line 20). Last but not least, all the selected test entities $\Gamma_{A}$, $\Gamma_{M}$, $\Gamma_{C}$, $\delta_{T}$ are used to rewrite their corresponding test classes to make sure only the selected assertions, methods or classes will be executed (Line 25). After the execution is finished, the test classes are restored to the original status, while the rewritten version is saved as a reference.
@@ IV Evaluation
  • RQ4 (effectiveness): How effective are the selected test cases in detecting faults?
  • RQ5 (performance): What is the performance of Selertion in terms of analysis overhead and collection time? Is there any performance difference when disabling the assertion-level selection?
@@ IV-A Existing RTS Tools
Our search criteria for existing test selection techniques to compare with included tools that are available, runnable, and target Java/JUnit. We selected two state-of-the-art tools, namely HyRTS [12] and Ekstazi [9]. To the best of our knowledge, these are the exclusive options that currently meet our criteria. Significantly, they outperform older RTS techniques like FaultTracer [11], which exhibit slower performance and offer test method selection, as documented in prior research [9].
HyRTS conducts the analysis of test dependencies at either the method or class level and selects test entities at the test class level. Ekstazi analyzes production code to identify changes at the class level. Although the original paper [9] mentions the provision of both method and class-level selection, the Ekstazi tool [23] is currently available with default settings that primarily select tests at the test class level. Our technique collects test dependencies also at a mix of method and class level similar to HyRTS, however, it has the ability to select test entities at multiple levels from the coarse-grained class to the fine-grained assertion level based on its analysis of the test suite. Therefore, theoretically speaking, Selertion should be able to select tests more precisely than both HyRTS and Ekstazi. We empirically assess this in our evaluation.
@@ IV-B Subject Systems
We include a total of 11 subject systems. These subjects are selected based on their relevance to prior research on test assertion analysis [34, 14] and regression test selection [12] and their popularity on GitHub. To ensure fair comparison among Selertion, HyRTS and Ekstazi, our selection criteria include subjects that use Maven with executable and passing JUnit4 test cases. We focus on subjects with a test runtime of at least 30 seconds. As reported by others [12, 9], for lower test runtimes it can take more time to analyze the code than to execute the entire test suite.
The selected subjects and their characteristics are shown in Table I with the ascending order of their testing time. All the subjects are single-module projects, except Tika [30] and Accumulo [31]. Because some modules of these multi-module projects only have a few test cases, for simplicity, we choose the module with most test cases to analyze (i.e., tika-parsers module for Tika, core module for Accumulo). We examine the test suites for each subject and blacklist test classes that do not meet the selection criteria. Specifically, since our implementation focuses on JUnit4 assertions and identifies test cases with the @Test annotation, we blacklist test classes that mix JUnit3 and JUnit4 APIs for the OpenTripPlanner subject. For the tika-parsers subject, we choose to skip test classes that mix Mockito [35] and JUnit4 APIs. Additionally, 9 out of the 11 subject systems, except for tabula-java and stream-lib, have recently migrated from JUnit4 to JUnit5 [36]. To maintain consistency, we start from a revision that uses JUnit4 as the Head to collect 100 revisions before it, and remove revisions that fail to build or have no code changes. For the three subjects commons-exec, commons-net and commons-pool, which lack a sufficient number of revisions with substantial source code changes, we introduce changes artificially. Because Selertion detects production code changes at a coarse-grained class or method level, our focus is on the location of the changes rather than the specifics of the changes themselves. By leveraging generated mutants, we can introduce code changes at various locations within the production code. We believe that these mutants can effectively mimic real code changes in this context. To achieve this, we use a source code mutation generator called Major [37] to create random mutants in various production methods, thus introducing more prevalent source code changes. The selected subjects span different application domains and exhibit diverse code and test suite sizes. Table I presents the short hash under the Head column, and lines of code for both production and test code for the Head revision. We measure lines of code using Cloc [38]. The column Time details the average time required to execute all the tests using the build command (i.e., mvn test) for the revisions we analyzed, along with the average execution time for each test case. This table also presents information pertaining to test cases and test assertions that are further elaborated in the following subsection.
@@ IV-C Procedure and Results
Given that we exclude parameterized tests, inherited tests, and tests dependent on other tests from our assertion-level selection (such tests are selected at method or class level), with consideration for the safety of our analysis, it is important to know to what extent the test suite for each subject can be fine-grained selected at the assertion-level. To measure the prevalence of fine-grained selection, we implemented a tool using JDT [19]. This tool statically counts the total number of test cases that can be fine-grained selected for every revision we executed. We present the average values for each subject in Table I. Additionally, because the smallest unit of interest for our study is the test assertion, we count the total number of assertions and calculate the average number per test. These numbers are included under the Assertions column in the table, both in total and per test. It’s worth noting that we opt to count these numbers statically instead of relying on information obtained from executed tests (e.g., as provided by maven-surefire-plugin [39]) to avoid repetitive counting, as a single test case may be executed multiple times due to inheritance or different parameters.
The column Fine-grained in Table I shows the number and percentage of test cases that can be fine-grained selected by Selertion in the subject systems. The results show on average 48% out of the total 8,328 test cases can be selected at the assertion level. The percentages range from 11% (commons-pool) to 75% (asterisk-java). Moreover, if each test case contains less than two assertions, the benefits of fine-grained selection become limited because the assertion-level selection will be equivalent to the method-level selection (i.e., the whole test method is selected). Therefore, we count explicit assertion statements (e.g., assertTrue ) inside each test method body. Column Assertions in Table I shows the number of assertions for each subject, which totals 28,177 for all subjects. The average number of assertions per test case ranges from 2.1 for commons-exec to 4.8 for tika-parsers, and the overall average of all the subjects is 3.3 assertions per test case. This indicates that (1) most test cases contain multiple assertions in practice, (2) tests that are analyzable by our approach at the fine-grain assertion-level are prevalent, and (3) our fine-grained selection approach is feasible to be applied to real-world test suites.
@@
To assess the efficacy of our approach in terms of precision, we use the widely used RTS metric Selected Test Ratio [40, 9, 41, 42, 10, 11, 12] to compare with the other two RTS tools. Selected Test Ratio is defined as the ratio of the number of selected test cases to the number of total test cases. As the fine-grained selection is at the test assertion level in this study, we define a new metric Selected Assertion Ratio as the ratio of the number of selected assertions to the number of total assertions. Lower selected test/assertion ratio indicates better selection precision, as it signifies the selection of fewer tests.
@@
Our precision evaluation results are presented in Table II and Figure 4. In Table II, the macro-columns Selected Assertion Ratio and Selected Test Ratio present the mean values of all the revisions we studied for Selertion, HyRTS and Ekstazi, respectively, except for commons-exec, tabula-java and tika-parsers. In the case of these subjects, HyRTS failed to run due to StackOverflow errors. Therefore, we mark the failed subjects and the corresponding average values for HyRTS with an asterisk to indicate that these values are calculated by excluding the subjects that failed, resulting in an average calculation different from that of Ekstazi. For a fair comparison, we calculate two average values for Selertion: one is the average across all subjects for comparison with Ekstazi, and the other, marked with an asterisk, is the average value excluding the HyRTS-failed subjects for comparison with HyRTS. As our results show, Selertion has lower selected test ratio values, which means it selects fewer test cases and assertions than the other two tools. Regarding selected assertion ratio, on average, Selertion selects 15.10% out of all the test assertions and all subjects, which is 52.1% less than Ekstazi. When compared with HyRTS (excluding the marked subjects), the average of Selertion is 10.63% which is 40.8% less than HyRTS’s average of 17.96%. For selected test ratio, overall our tool selects 15.80% of all the test cases (10.85% when excluding the marked subjects). This is 43.4% less than HyRTS and 53.0% less than Ekstazi. For each subject, the precision increase in terms of selected assertion ratio is from 10.8% (OpenTripPlanner) to 65.1% (commons-math) in comparison with HyRTS, and from 16.1% (tabula-java) to 80.8% (commons-math) compared with Ekstazi. Meanwhile, our tool decreases the selected test ratio from 22.9% (commons-pool) to 94.4% (commons-net) in comparison to HyRTS, and from 18.2% (tabula-java) to 94.9% (commons-net) compared with Ekstazi. It is worth mentioning that the selected ratios are related to the size of code changes in each sample revision we collect to analyze. Larger code changes always lead to higher selected ratios. Moreover, to provide more insights about selected assertion ratios for each subject besides the mean values, Figure 4 includes boxplots depicting the distributions of selected assertion ratios for each revision. The mean values are marked as diamonds, and median values are marked as lines. The results indicate that our approach has the ability to select tests more precisely than other coarser-grained tools.
@@
Because our approach is implemented in an offline fashion, where the data collection phase is executed in a separate process and the instrumented source code is stored in a separate copy of the original project, we measure the offline End-to-End time (AE time) for Selertion, HyRTS, and Ekstazi, respectively, to ensure a fair comparison and assess the efficiency in reducing testing time. The offline End-to-End time for Selertion includes the analysis (A) time to compute code changes, select affected test entities and update assertion slices, as well as the execution (E) time for the selected test entities, then normalize the end-to-end time into a ratio with respect to the overall time of executing the whole test suite (RetestAll). We perform the measurement for each revision and report the average. We also measure the offline end-to-end time and its ratio for HyRTS and Ekstazi respectively. The results are shown in Table III. Column RetestAll shows the time required to run the whole test suite in average for each subject as a reference. The other three macro-columns present the offline end-to-end time and its ratio to RetestAll for Selertion, HyRTS, and Ekstazi separately. Besides the average values in Table III, to further understand the results, we perform Paired T-Test [43] with the significance level as 5% ($\alpha$ = 0.05) to determine if there is a significant difference between each two groups (i.e., Selertion vs. HyRTS, Selertion vs. Ekstazi and HyRTS vs. Ekstazi) for all the revisions of every subject. The reason to choose Paired T-Test is that as a parametric test, it has greater statistical power than non-parametric tests like Wilcoxon Test [43], and can provide trustworthy results even for non-normal distributions. If the end-to-end time has no statistical difference in comparison with others, it means it has competitive or equivalent capability for reducing test time. After examining the results of offline end-to-end time, in Table III, we mark the subject as light gray if Selertion performs competitively or better than one of the other tools; otherwise, Selertion outperforms the other two tools for that subject. We also highlight the shortest average running time in yellow in each row across the three tools.
Across all 11 subjects, Selertion consistently reduces testing time effectively, with ratios to RetestAll ranging from 14.9% (commons-net) to 70.3% (OpenTripPlanner). This reduction translates to an overall decrease in testing time of 29.7% to 85.1%. For asterisk-java with 35 seconds runtime, our tool is competitive with HyRTS, but slower than Ekstazi. For subject commons-net, although HyRTS has the shortest running time, Selertion is only 1 second slower than HyRTS and is equivalent to Ekstazi, which has no statistical difference in comparison with both tools. For subject commons-math, Selertion outperforms Ekstazi but is 5.3 seconds slower than HyRTS.
@@
Longer Test Runtime. For the subjects with runtimes longer than 86 seconds, Selertion performs significantly better than Ekstazi in terms of testing time. HyRTS fails to run on the subjects commons-exec, tabula-java and tika-parsers, thus we treat Selertion as superior on these three subjects. As an exception, commons-math has long test time with 160 seconds, the reason that HyRTS works better than Selertion is the execution time per test for commons-math is very fast with only 0.03s (shown in Table I). Therefore, although test assertions are reduced significantly, the reduced execution time is marginal. Although HyRTS has lower average testing time for commons-math (row 8), the result of Paired T-Test indicates that there is no significant difference between Selertion and HyRTS. Selertion performs better than HyRTS for the remaining subjects. Generally speaking, we can conclude that Ekstazi works best for subjects with short test runtime such as asterisk-java with 35s running time, Selertion outperforms or is competitive to both HyRTS and Ekstazi with long-running subjects i.e., test runtime longer than 86 seconds in our experiments. The results show that our fine-grained RTS technique is more effective in handling time-intensive tests.
Considering all test suites in our subjects, short and long, the total end-to-end time on average for Selertion is 74.33 seconds while the average overall test execution time (RetestAll) is 200.36 seconds. To the ratio of RetestAll, our technique reduces 63% of the overall test execution time. Compared with Ekstazi which has an average end-to-end time of 119.74 seconds, Selertion is overall 38% faster than Ekstazi. Compared with HyRTS when we ignore commons-exec, tabula-java and tika-parsers, the average end-to-end time for HyRTS is 83.51 seconds and 77.51 seconds for Selertion which is 7% faster than HyRTS. The average gain compared with HyRTS could be potentially larger if HyRTS were executable on commons-exec, tabula-java and tika-parsers. Our results show that Selertion is more precise and more efficient, especially for test suites with longer execution times, which benefit most from test selection in practice.
@@
Ave. 200.36 74.33 / 77.51* 41.49 / 37.25* 83.51* 38.40* 119.74 54.33
TABLE IV: Killed mutants by the original test suite and RTS tools
Subject (Original) Selertion HyRTS Ekstazi
asterisk-java 2 2 2 2
commons-net 3 3 3 3
commons-exec* 2 2 - 2
tabula-java* 3 3 - 3
OpenTripPlanner 1 1 1 1
commons-math 5 5 5 5
stream-lib 5 5 5 5
tika-parsers* 3 3 - 3
accumulo-core 5 5 5 5
commons-pool 2 2 2 2
LogicNG 5 5 5 5
IV-C4 Effectiveness (RQ4).
A safe RTS tool should ensure that the number of newly failed tests when run with the tool matches the number when running the original test suite [44]. To asset the effectiveness of Selertion and compare it with state-of-the-art approaches, we employ Major [37] to randomly generate ten mutants across the production code for each subject and then count the number of eliminated mutants for the original test suite, Selertion, HyRTS and Ekstazi. The results are reported in Table IV. For all subjects, the total number of mutants killed, as well as each individual mutant eliminated by the RTS tools Selertion, HyRTS and Ekstzai, align with those killed by the original test suite. These empirical findings affirm that our fine-grained selection approach maintains the same test effectiveness as observed in the original test suite, demonstrating that it is safe. Additionally, we discuss other efforts taken to ensure the safety of Selertion in Section V.
IV-C5 Performance (RQ5).
TABLE V: Performance results for Selertion and comparison to a method-level variant